The product of 2 consecutive positive even integers is 360. Find the integers.
Answer by checkley75(3666)

X(X+2)=360
X^2+2X-360=0
(X-18)(X+20)=0
X-18=0
X=18 ANSWER.
X+20=0
X=-20 ANSWER.
PROOF
18(18+2)=360
18*20=360
360=360